OpenStudy (love_333):

Medal + fan for answer & work out of this problem: Find the area of a regular hexagon with an apothem 10.4 yds long and a side 12 yds long. Round your answer to the nearest tenth.

OpenStudy (love_333):

@563blackghost Could you help me one more time to make sure i've got this down.

563blackghost (563blackghost):

Use the equation.... \(\Huge{\frac{1}{2}(a \times p)}\) We are aware of the apothem but we need to know of the perimeter so we would multiply `12` by the number of sides....

OpenStudy (love_333):

So, 72 correct?

563blackghost (563blackghost):

Correct :) Now we input the apothem and the perimeter in the equation... \(\Huge{A=\frac{1}{2}(10.4 \times 72)}\)

OpenStudy (love_333):

374.4?

563blackghost (563blackghost):

Yup :)

OpenStudy (love_333):

Yay!! Thank you!

563blackghost (563blackghost):

np ^^